# An Analytical and AI-discovered Stable, Accurate, and Generalizable Subgrid-scale Closure for Geophysical Turbulence

### v2

@@ Acknowledgment
This work is supported by NSF grant AGS-2046309 and Schmidt Sciences. Computational resources were provided by NSF ACCESS (allocation ATM170020) and NCAR’s CISL (allocation URIC0009).
@@
Figure 4: The four cases considered in this study. These cases represent a broad range of dynamics in 2D turbulence and mimic the diversity of flow regimes (jets and vortices) in the atmosphere and ocean. $(k_{x},k_{y})$ are the wavenumbers of the time-invariant forcing in the $x$ and $y$ directions and $\beta$ is the gradient of the Coriolis force (see [21] for details). $N_{\mathrm{DNS}}$ is the resolution of the pseudo-spectral solver (in each direction) used for DNS (Jakhar et al. [71]). The first row presents snapshots of the DNS vorticity, $\omega$. The second row shows snapshots of the FDNS vorticity, $\overline{\omega}$. The third row displays a snapshot of an element of the subgrid-scale (SGS) term, $\tau_{12}$. The last row depicts the energy (red line) and enstrophy (blue line) spectra for DNS (dashed line) and FDNS (solid line). FDNS is at $N_{\text{LES}}=128$ for Cases 1-2 and $N_{\text{LES}}=512$ for Cases 3-4.
Appendix B: The baseline physics-based closures used here are Smag, DSmag, and DLeith. These eddy viscosity closures introduce dissipation and do not account for backscattering: $\tau_{ij}=-2\nu_{e}\overline{S}_{ij}$, with $\overline{S}_{ij}$ representing the filtered rate of strain and $\nu_{e}$ denoting the eddy viscosity [1]. $\nu_{e}=\left(C_{s}\Delta\right)^{2}\sqrt{2\overline{S}_{ij}\overline{S}_{ij}}$ for Smag and $\nu_{e}=\left(C_{l}\Delta\right)^{3}|\nabla\overline{\omega}|$ for Leith. We use $C_{s}=0.17$ for Smag, as proposed in a previous work [22]; $C_{s}$ and $C_{l}$ are estimated dynamically based on the local flow structure for DSmag and DLeith, respectively. This procedure can yield $\nu_{e}\leq 0$, potentially resulting in numerical instabilities; therefore, the commonly used “positive clipping” is applied to enforce diffusion ($\nu_{e}\geq 0$) [74].

### v3

@@ Acknowledgment
We thank two anonymous reviewers for insightful feedback. This work is supported by NSF grant AGS-2046309 and Schmidt Sciences. Computational resources were provided by NSF ACCESS (allocation ATM170020) and NCAR’s CISL (allocation URIC0009).
@@
Author Contributions
All authors participated in conceptualizing the project, designing the computational experiments, and interpreting the results. K.J. developed the codes, performed the analysis, and generated the figures. P.H. acquired the funding and supervised the research.
@@
Figure 4: The four cases considered in this study. These cases represent a broad range of dynamics in 2D turbulence and mimic the diversity of flow regimes (jets and vortices) in the atmosphere and ocean. $(k_{x},k_{y})$ are the wavenumbers of the time-invariant forcing in the $x$ and $y$ directions, where higher wavenumbers lead to small-scale vortices and double cascade. $\beta$ is the gradient of the Coriolis force, mimicking the effects of Earth’s rotation and leading to the formation of jets (see [21] for more details about the dynamics of these cases). $N_{\mathrm{DNS}}$ is the resolution of the pseudo-spectral solver (in each direction) used for DNS (Jakhar et al. [74]). The first row presents snapshots of the DNS vorticity, $\omega$. The second row shows snapshots of the FDNS vorticity, $\overline{\omega}$. The third row displays a snapshot of an element of the SGS stress term, $\tau_{12}$ (the subject of closure modeling). The last row depicts the energy (red line) and enstrophy (blue line) spectra for DNS (dashed line) and FDNS (solid line). FDNS is at $N_{\text{LES}}=128$ for Cases 1-2 and $N_{\text{LES}}=512$ for Cases 3-4.
Appendix B: Correlation coefficienct (CC) compares the 2D pattern of the “true” SGS stress, $\tau^{\text{FDNS}}$ (from FDNS), with the predicted stress, $\tau^{\text{RVM}}$, which is generated by the RVM-discovered closure based on the filtered flow variables (e.g., $\overline{u}_{1}$, $\overline{u}_{2}$) as input:
$\text{CC}=\frac{\langle\left(\tau^{\text{RVM}}-\langle\tau^{\text{RVM}}\rangle\right)\left(\tau^{\text{FDNS}}-\langle\tau^{\text{FDNS}}\rangle\right)\rangle}{\sqrt{\langle\left(\tau^{\text{RVM}}-\langle\tau^{\text{RVM}}\rangle\right)^{2}\rangle}\sqrt{\langle\left(\tau^{\text{FDNS}}-\langle\tau^{\text{FDNS}}\rangle\right)^{2}\rangle}},$ (5)
where $\langle\cdot\rangle$ is domain averaging.
Appendix C: The DNS/LES numerical solver is the same as the one used in [37, 21]. Briefly, the Navier-Stokes equations in the vorticity-streamfunction formulation, e.g., curl of Eq. (2) for LES, are solved in a doubly periodic domain using a Fourier-Fourier pseudo-spectral solver. For time integration, we employ second-order Adams-Bashforth and Crank-Nicolson schemes for the advection and viscous terms, respectively.
Appendix D: The baseline physics-based closures used here are Smag, DSmag, and DLeith. These eddy viscosity closures introduce dissipation and do not account for backscattering: $\tau_{ij}=-2\nu_{e}\overline{S}_{ij}$, with $\overline{S}_{ij}$ representing the filtered rate of strain and $\nu_{e}$ denoting the eddy viscosity [1]. $\nu_{e}=\left(C_{s}\Delta\right)^{2}\sqrt{2\overline{S}_{ij}\overline{S}_{ij}}$ for Smag and $\nu_{e}=\left(C_{l}\Delta\right)^{3}|\nabla\overline{\omega}|$ for Leith. We use $C_{s}=0.17$ for Smag, as proposed in a previous work [22]; $C_{s}$ and $C_{l}$ are estimated dynamically based on the local flow structure for DSmag and DLeith, respectively. This procedure can yield $\nu_{e}\leq 0$, potentially resulting in numerical instabilities; therefore, the commonly used “positive clipping” is applied to enforce diffusion ($\nu_{e}\geq 0$) [71].